OpenStudy (anonymous):

Can someone explain to me why this function, x=f^-1(y), then y=f(y). is false?

OpenStudy (loser66):

because y = f(x) is true so that y = f(y) must be false. hihihi

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

There is exception though. if y=x, then x=f^-1(y) is true, but y=f(y) is also true, because f(y)=y, hence y=f(y)=y

OpenStudy (anonymous):

the answer is false but i'm not sure why